OpenStudy (anonymous):

rewrite ln(2/3) in terms of ln4 and ln3

OpenStudy (slaaibak):

\[\ln({2 \over 3}) = \ln 2 - \ln 3\] but ln 4 = 2ln2 so ln 2 = (ln4)/2 so \[\ln({2 \over 3}) = {\ln 4 \over 2} - \ln 3\]

OpenStudy (anonymous):

thanks!